OpenStudy (anonymous):

Solve simultaneously for "x" and "y": x+2y=9 x-y=3

OpenStudy (bahrom7893):

x+2y = 9 x=3+y 3+y+2y = 9 x = 3+y

OpenStudy (bahrom7893):

3+3y=9 x=3+y 3y=6 x=3+y y=2 x=3+2=5